# On the Challenges and Opportunities in Generative AI

### v3

@@ 4.5 Constraint Satisfaction
Generative models such as ChatGPT are used by millions of people and deployed across diverse use cases. For many of these, they are expected to satisfy additional constraints. [^9] Here we focus on constraints specified at inference time. We discuss constraints that must be integrated into the model during training in 2.2. In some cases, these merely stem from a desire to have a more controlled form of generation, such as when a generated image is conditioned on a given depth map (Zhang et al., 2023a). In other cases, ethical and safety considerations are key concerns. For instance, in fields like engineering design, generative model outputs must meet engineering standards and adhere to laws of nature (i.e., physics). More generally, there are widespread calls for generative models to avoid toxicity, bias mitigation and other outputs that may lead to harmful effects (Weidinger et al., 2021), e.g., by refraining from responding in ways that could pose a risk to the mental health of human interlocutors and by refusing to carry out tasks related to illegal activities. While reinforcement learning from human feedback (Ouyang et al., 2022) has offered an initial step towards these goals, enabling companies and users to provide models with soft constraints within their queries, such constraints can be circumvented (Shen et al., 2023). Ultimately, methods that allow us to place hard constraints on model outputs are necessary.
@@
In language generation, decoding methods that allow for arbitrary constraints (both hard and soft) have been a focus area of the research community (Kumar et al., 2021; 2022; Beurer-Kellner et al., 2024). There are several challenges that proposed approaches, including: efficiently enforcing constraints without significantly increasing computational costs, maintaining fluency and coherence while adhering to constraints, and handling multiple constraints, which may result in conflicting requirements. The discrete nature of text presents a particular difficulty, as small changes to token sequences can drastically alter meaning, making it difficult to optimize for constraints from a computational perspective. Meanwhile, research on controllable image generation has likewise gained momentum (Deng et al., 2020; Huang et al., 2024b), with various approaches aiming to regulate attributes such as style, composition, or specific content elements. Methods range from applying spatial constraints (e.g., bounding boxes, masks, or layout specifications (Zheng et al., 2023a)) to enforcing semantic conditions (e.g., ensuring that certain objects or visual features are present (Pavllo et al., 2020)). Technical challenges arise in this domain as well, such as the need for more complex conditioning mechanisms and heavier computational demands—especially when constraints must be integrated at each step of the generative process in e.g., diffusion models. Code generation stands out as a domain where constraint enforcement has long been a primary focus (Poesia et al., 2022; Dong et al., 2023). Here, the constraints—such as following a language’s syntax and producing compilable code—are arguably more well-defined and straightforward to verify. Concepts and methods from this field could conceivably help research in enforcement of hard constraints in other generative AI fields. Overall, methods to ensure effective constraint adherence can substantially improve control over generative models, which is crucial for ensuring their safe and reliable deployment (Regenwetter et al., 2024). Despite progress across various generative AI fields, the development of scalable and generalizable techniques capable of handling the diverse and often conflicting demands of real-world constraints remains an open challenge. We encourage further research in this area to bridge this gap and advance the controllability of generative models across different domains.
@@ 5 Conclusion
Despite the recent excitement and hype surrounding advancements in generative AI, the goal of achieving a perfect generative model remains far from reality. In the hopes of eventually reaching this goal, we identified several core challenges with the current generative modeling frameworks and practices.
We identified generalization and robustness as major hurdles, demanding methods to handle unseen data and adversarial attacks. Moreover, limited representational power and implicit modeling and data assumptions necessitate exploring more expressive models and incorporating prior knowledge, especially in data-scarce scenarios. Moving beyond merely identifying correlation, integrating causal reasoning into DGMs holds promise for enhanced interpretability and robustness. We also emphasized the escalating computational demands and associated barriers to the widespread adoption of DGMs. Training and inference inefficiencies pose key challenges, urging the exploration of alternative network architectures and low-bitrate model quantization. Inadequate evaluation metrics for generated content, such as FID and n-gram matching, hinder efficient progress, prompting the search for robust, domain-agnostic alternatives.
We also discussed the challenges and considerations surrounding the responsible deployment of large-scale generative models, pointing to the rising concerns related to misinformation, unintended harm, and lack of trustworthiness. Challenges include combating misinformation, ensuring privacy in data curation, addressing fairness issues, enhancing interpretability, estimating model uncertainty, and satisfying constraints.
By confronting the limitations discussed here, we can transform DGMs from data replicators (Bender et al., 2021) to tools with transformative capabilities across various domains. We hope that our paper will point to directions that ultimately contribute to these goals.

### v4

@@ 4.5 Constraint Satisfaction
Generative models such as ChatGPT are used by millions of people and deployed across diverse use cases. Many applications require generative models to satisfy domain-specific constraints. [^10] Here we focus on constraints specified at inference time. We discuss constraints that must be integrated into the model during training in 2.3. In some cases, these merely stem from a desire to have a more controlled form of generation, such as when a generated image is conditioned on a given depth map (Zhang et al., 2023b). In other cases, ethical and safety considerations are key concerns. For instance, in fields like engineering design, generative model outputs must meet engineering standards and adhere to laws of nature (i.e., physics). More generally, there are widespread calls for generative models to avoid toxicity, mitigate bias and prevent other outputs that may lead to harmful effects (Weidinger et al., 2021), e.g., by refraining from responding in ways that could pose a risk to the mental health of human interlocutors and by refusing to carry out tasks related to illegal activities. While reinforcement learning from human feedback (RLHF; Ouyang et al., 2022)—and particularly constrained RLHF (Moskovitz et al., 2024)—has offered an initial step towards these goals, enabling companies and users to provide models with soft constraints within their queries, such constraints can be circumvented (Shen et al., 2023). Ultimately, methods that allow us to place hard constraints on model outputs are necessary.
@@
In language generation, decoding methods that allow for arbitrary constraints (both hard and soft) have been a focus area of the research community (Kumar et al., 2021; 2022). There are several prominent challenges in the development of such methods, including: efficiently enforcing constraints without significantly increasing computational costs, maintaining fluency and coherence while adhering to constraints, and handling multiple constraints, which may result in conflicting requirements. The discrete nature of text presents a particular difficulty, as small changes to token sequences can drastically alter meaning, making it difficult to optimize for constraints from a computational perspective. Recent grammar-constrained decoding methods (Beurer-Kellner et al., 2024; Ugare et al., 2024) address these issues by guaranteeing that every generated sequence conforms to a user-supplied context-free grammar, achieving hard constraints with only a modest runtime overhead. Such methods have started to gain traction in other domains, e.g., in healthcare (Golechha & Dao, 2024b). Meanwhile, research on controllable image generation has likewise gained momentum (Deng et al., 2020; Huang et al., 2024b), with various approaches aiming to regulate attributes such as style, composition, or specific content elements. Methods range from applying spatial constraints (e.g., bounding boxes, masks, or layout specifications (Zheng et al., 2023a)) to enforcing semantic conditions (e.g., ensuring that certain objects or visual features are present (Pavllo et al., 2020)). Technical challenges arise in this domain as well, such as the need for more complex conditioning mechanisms and heavier computational demands—especially when constraints must be integrated at each step of the generative process in e.g., diffusion models. Code generation stands out as a domain where constraint enforcement has long been a primary focus (Poesia et al., 2022; Dong et al., 2023). Here, the constraints—such as following a language’s syntax and producing compilable code—are arguably more well-defined and straightforward to verify. Concepts and methods from this field could conceivably help research in enforcement of hard constraints in other generative AI fields, e.g., in the application of generative AI to the physical domain, where laws of nature must be satisfied. Overall, methods to ensure effective constraint adherence can substantially improve control over generative models, which is crucial for ensuring their safe and reliable deployment (Regenwetter et al., 2024). Despite progress across various generative AI fields, the development of scalable and generalizable techniques capable of handling the diverse and often conflicting demands of real-world constraints remains an open challenge. We encourage further research in this area to bridge this gap and advance the controllability of generative models across different domains.
@@ 5 Conclusion
Generative AI has achieved remarkable progress in recent years, pushing the boundaries of what computational systems can model and generate. Yet, this rapid evolution also comes with a wide array of technical, ethical, and societal challenges that demand careful attention. Thus, the goal of achieving a perfect generative model remains far from reality. As we have outlined throughout this paper, current generative AI models struggle with robustness under distribution shift, remain resource-intensive, and often exhibit behaviors that are difficult to interpret or control, particularly in high-stakes and specialized domains.
Crucially, the path forward is not merely a matter of scaling existing paradigms. Despite gains in performance from ever-larger models and training corpora, core limitations persist. Addressing them requires (re)consideration of foundational assumptions. We have emphasized the importance of integrating causal reasoning into model architectures to overcome spurious correlations and improve generalization, especially under distributional shifts. Likewise, embedding domain knowledge and inductive biases can enable more data-efficient learning in fields like healthcare, chemistry, and the physical sciences, where data is often limited but expert insight is abundant.
Efficiency and accessibility are also central concerns for facilitating the widespread use of DGMs. Today’s DGMs are not only expensive to train and deploy but also have created a growing barrier to entry for researchers and practitioners without access to large-scale compute. Optimizing inference efficiency through quantization, architectural innovations, and fast sampling methods is essential for democratizing generative AI and minimizing its environmental footprint. At the same time, robust evaluation remains an open challenge: current metrics often fail to align with human judgments or to meaningfully assess qualities such as diversity, calibration, or ethical compliance. The development of more principled and context-dependent evaluation frameworks—potentially learned from human preferences—will be vital to guiding model development and ensuring model trustworthiness.
Beyond technical concerns, we must also address the broader societal and ethical impacts that generative AI can have. Interpretability, fairness, and safety must become a priority in design considerations, not an afterthought. These systems have already demonstrated the ability to generate misinformation, propagate social biases, infringe on privacy, and produce unsafe or unreliable outputs. While recently-developed mitigation strategies, from watermarking and uncertainty-aware abstention to differential privacy and constraint-aware decoding, appear to offer promising solutions, the methods alone are insufficient without regulatory frameworks for ensuring their usage.
We believe that addressing this range of issues will require deeper cross-disciplinary collaboration, new benchmarks that reflect real-world complexity, and an intentional shift from optimizing for performance to optimizing for positive societal impact. By confronting the limitations discussed here, we can transform DGMs from data replicators (Bender et al., 2021) to tools with transformative capabilities across various domains. We hope that the information and opinions presented here will point to directions that ultimately contribute to these goals.